# Lee-Yang-zero ratios for locating a critical point

### v1

@@ Abstract
In the present Letter, we focus on the use of the FSS of Lee-Yang zeros (LYZ) [17, 18], i.e. zeros of the partition function on the complex-variable space, for numerical investigations of a CP in general systems. Whereas the FSS of LYZ has been investigated in the literature [19, 20, 21, 22, 23], to the best of our knowledge its systematic utilization for this purpose has not been discussed so far. The LYZ have been investigated in lattice-QCD numerical simulations [24, 25, 26]. In particular, its application for locating the CP in QCD at nonzero chemical potential has been discussed recently [27, 28, 29, 30] in connection to the Lee-Yang edge singularity (LYES) [31, 32, 33, 34, 35, 36, 37, 38]. However, the finite-volume effects have not been investigated in detail in these studies.
@@
As a systematic procedure to study a CP in numerical simulations, based on the FSS we show that the ratios of the imaginary parts of LYZ obtained on different volumes intersect at the CP in the large-volume limit. We propose the use of this property, which is similar to the Binder cumulant [14], for locating a CP in numerical simulations. The LYZ carry information of the system that is not encoded in finite derivatives of the free energy. Our method thus serves as a novel procedure independent of the ordinary methods that rely on susceptibilities. We test the method for analyzing the CP in the three-dimensional three-state Potts model and show that the method can determine its location successfully with almost the same statistical error compared to the Binder-cumulant method. As byproducts, some useful properties of LYZ, especially their relation to the LYES, will also be clarified.
@@
Ising model.— To illustrate the method, we start from a simple case of the conventional three-dimensional Ising (3d-Ising) model [3] described by the reduced temperature $t$ and external magnetic field $h$ on the cubic lattice of size $L^{3}$, having a CP at $(t,h)=(0,0)$ and a first-order phase transition at $h=0$ for $t<0$ for $L\to\infty$. We denote the partition function of this model as $Z(t,h,L^{-1})$, which satisfies $Z(t,h,L^{-1})=Z(t,-h,L^{-1})$. The LYZ of this model are the values of $h\in\mathbb{C}$ satisfying $Z(t,h,L^{-1})=0$ for a given $t\in\mathbb{R}$ [17, 18]. It is known as the Lee-Yang circle theorem that the LYZ distribute discretely on the pure-imaginary axis for finite $L$ [18]. In the following, we denote the LYZ with ${\rm Im}h>0$ as $h=h_{\rm LY}^{(n)}(t,L)$, where $n=1,2,\cdots$ labels different LYZ so that $0<{\rm Im}h_{\rm LY}^{(1)}(t,L)<{\rm Im}h_{\rm LY}^{(2)}(t,L)<\cdots$; the system has the other LYZ at $h=-h_{\rm LY}^{(n)}(t,L)$. Since $Z(t,h,L^{-1})$ at finite $L$ is a regular function of $t$ and $h$, $h_{\rm LY}^{(n)}(t,L)$ should also be a regular function of $t$ for finite $L$.
@@
for sufficiently large $L$. In the 3d-Ising model, the values of the exponents, $y_{t}\simeq 1.588$ and $y_{h}\simeq 2.482$, have been analyzed with high precision [39, 15, 16]. Since the LYZ are given by zeros of Eq. (1), it immediately follows that the LYZ for different $L$ obey
@@
for finite $n$. Since the right-hand side of Eq. (3) does not depend on $L$, only a possible asymptotic behavior of ${\tilde{h}}_{\rm LY}^{(n)}({\tilde{t}})$ for $\tilde{t}\to\infty$ is ${\tilde{h}}_{\rm LY}^{(n)}({\tilde{t}})\propto{\tilde{t}}^{y_{h}/y_{t}}$, which yields $h_{\rm LYES}(t)\propto t^{y_{h}/y_{t}}$ [21].
@@
with $r_{nm}=X_{n}/X_{m}$ and $c_{nm}=r_{nm}(Y_{n}/X_{n}-Y_{m}/X_{m})$. Equation (6) shows that $R_{nm}(0,L)=r_{nm}$ is independent of $L$, while the slope at $t=0$ scales as $L^{y_{t}}$. In other words, $R_{nm}(t,L)$ for different $L$ intersect at $t=0$ as in Fig. 1 (a). This property is similar to the Binder cumulant [14] and would provide an alternative method to determine the critical temperature from the intersection point in numerical simulations.
@@
CP in general systems.— Next, we extend the argument to CPs in general systems that belong to the same universality class as the 3d-Ising model. We suppose a system whose partition function ${\cal Z}(\tau,\xi,l^{-1})$ is described by two variables $\tau$ and $\xi$ with $l$ being a dimensionless parameter proportional to the spatial size of the system. We also assume that this system has a first-order phase-transition line on the $\tau$$\xi$ plane that terminates at a CP at $(\tau,\xi)=(\tau_{\rm c},\xi_{\rm c})$ as depicted in Fig. 1 (b). In the following, we consider zero points of the partition function for $\xi\in\mathbb{C}$ and $\tau\in\mathbb{R}$ and refer to them as the LYZ $\xi=\pm\xi_{\rm LY}^{(n)}(\tau,l)$, i.e. ${\cal Z}(\tau,\pm\xi_{\rm LY}^{(n)}(\tau,l),l^{-1})=0$, where the definition of the label $n$ is the same as before.
Since the CP belongs to the 3d-Ising universality class and there are only two relevant variables at the CP, the partition function in the vicinity of the CP for large $l$ is related to that of the Ising model as
@@
$l^{y_{h}}\check{h}\big{(}\tau,\xi_{\rm LY}^{(n)}(\tau,l)\big{)}=\tilde{h}_{\rm LY}^{(n)}\big{(}l^{y_{t}}\check{t}(\tau,\xi_{\rm LY}^{(n)}(\tau,l))\big{)}.$ (10)
@@
$l^{y_{h}}\big{(}a_{21}\delta\tau+a_{22}(\xi_{\rm LY}^{(n)}(\tau,l)-\xi_{\rm c})\big{)}$
$=i\big{(}X_{n}+Y_{n}l^{y_{t}}\big{(}a_{11}\delta\tau+a_{12}(\xi_{\rm LY}^{(n)}(\tau,l)-\xi_{\rm c})\big{)}\big{)}+{\cal O}(\delta\tau^{2}),$ (11)
@@
Now, let us compare Eq. (15) with the Binder cumulant [14]. For locating a CP on the $\tau$$\xi$ plane, one may define the Binder cumulant as ${\cal B}_{4}(\tau,l)={\rm min}_{\xi}[(\partial^{4}{\cal F}(\tau,\xi,l^{-1})/\partial\xi^{4})/(\partial^{2}{\cal F}(\tau,\xi,l^{-1})/\partial\xi^{2})^{2}]+3$ [41, 42, 43] with the free energy ${\cal F}(\tau,\xi,l^{-1})=-T\ln{\cal Z}(\tau,\xi,l^{-1})$ ) with temperature $T$. One then obtains [42, 44]
@@
${\cal H}(\tau,\xi)=-\tau\sum_{\langle i,j\rangle}\delta_{\sigma_{i}\sigma_{j}}-\xi\sum_{i}\delta_{\sigma_{i},1},$ (17)
on the simple cubic lattice of size $L^{3}$ with periodic boundary conditions, where $\sigma_{i}$ takes three states $\sigma_{i}=1,2,3$ with the subscript denoting the lattice site and $\sum_{\langle i,j\rangle}$ represents the sum over all pairs of adjacent sites. We employ the notation that the partition function is given by ${\cal Z}={\rm Tr}e^{-{\cal H}(\tau,\xi)}$ following Ref. [41]. As schematically shown in Fig. 1 (b), this model has a first-order phase transition at vanishing external field $\xi=0$ owing to $Z(3)$ symmetry, which is eventually terminated at a CP for $\xi>0$ that belongs to the 3d-Ising universality class [41]. In Ref. [41], this CP has been investigated by the Binder-cumulant method in relation to the CP in QCD with heavy-mass quarks [44, 43, 45].
@@
with $\tau\in\mathbb{R}$ and $\xi\in\mathbb{C}$ for each simulation parameter, where $\langle\cdot\rangle_{L}$ denotes the average over the configurations at $(\tau_{\rm sim},\xi_{\rm sim})$ of size $L^{3}$. Whereas the analysis of Eq. (18) suffers from the overlapping problem when $(\tau,\xi)$ is largely deviated from $(\tau_{\rm sim},\xi_{\rm sim})$, we found that this problem is well suppressed in our analysis as demonstrated below.
@@
Table 1: Fit results of the CP parameters and $\chi^{2}/{\rm dof}$.
fit data $\tau_{\rm c}$ $y_{t}$ $r_{n1}$ or $b_{4}$ $\chi^{2}/\rm{dof}$
@@
To obtain the critical value $\tau=\tau_{\rm c}$ at the CP, we performed the four-parameter chi-square fits to the data of ${\cal R}_{n1}(\tau,L)$ for $L\geq 40$ ($12$ data points in total) with an ansatz ${\cal R}_{n1}(\tau,L)=r+c(\tau-\tau_{\rm c})L^{y_{t}}$ with $r$, $c$, $\tau_{\rm c}$, and $y_{t}$ being the fitting parameters. Effects of the second bracket in Eq. (15) are neglected since no deviation of the intersection point is visible for $L\geq 40$ in Fig. 2. The fit results are shown by the diamonds in Fig. 2 and summarized in Table 1. The table shows that the results for $n=2,3,4$ are consistent with each other.
@@
Summary and outlook.— In this Letter, we discussed the FSS of LYZ and showed that the intersection point of their ratios, Eq. (4) or Eq. (16), for various spatial volumes indicates a CP in general systems. This property can be used for numerical searches of CPs as an independent method from the conventional ones based on susceptibilities. Compared to the Binder-cumulant method [14], this method is advantageous in suppressing the finite-volume effects arising from $a_{12}\neq 0$. We applied the method to the numerical analysis of the CP in the three-state Potts model. Whereas we assumed the CP in the 3d-Ising universality class throughout the Letter for a simple presentation, this method, of course, can be adapted to CPs in other universality classes.
In our numerical study, we limited our analysis to the LYZ with $n\leq 4$. However, one can use the LYZ for yet larger $n$, which will act to improve the statistics. The LYZ can also be used to determine $\xi_{\rm c}$ and the matrix $A$, where better control of the finite-volume effects and statistics will be realized by combined use of Eqs. (13) and (14) for various $n$. The precise measurements of $r_{nm}$ in individual universality classes and other quantities [16], as well as the application of the method to CPs in various systems, such as the QCD critical point in lattice simulations, are other important future studies.
Acknowledgement.— We thank Shinji Ejiri for discussions in the early stage of this study. The authors also thank Koji Hukushima for useful discussions. This work was supported in part by JSPS KAKENHI (Nos. JP19H05598, JP22K03593, JP22K03619, JP23H04507, JP24K07049), the Center for Gravitational Physics and Quantum Information (CGPQI) at Yukawa Institute for Theoretical Physics, the Research proposal-based use at the Cybermedia Center, Osaka University, and the Multidisciplinary Cooperative Research Program of the Center for Computational Sciences, University of Tsukuba.

### v3

# Locating Critical Points Using Ratios of Lee-Yang Zeros

@@ Abstract
In the present Letter, we focus on the use of the FSS of Lee-Yang zeros (LYZ) [17, 18], i.e. zeros of the partition function on the complex-variable space, for numerical investigations of a CP in general systems. Whereas the FSS of LYZ has been investigated in the literature [19, 20, 21, 22, 23], to the best of our knowledge its systematic utilization for this purpose has not been discussed so far. The LYZ have been investigated in lattice-QCD numerical simulations [24, 25, 26, 27, 28]. In particular, its application for locating the CP in QCD at nonzero chemical potential has been discussed recently [29, 30, 31, 32] in connection to the Lee-Yang edge singularity (LYES) [33, 34, 35, 36, 37, 38, 39, 40]. However, the finite-volume effects have not been investigated in detail in these studies. As we discuss below, systematic utilization of the FSS of LYZ provides us with a general procedure applicable to a wide variety of numerical simulations including lattice QCD and those in statistical physics.
@@
We show that the ratios of the imaginary parts of LYZ obtained on different volumes intersect at the CP in the large-volume limit. We propose the use of this property, which is similar to the Binder cumulants [14], for locating a CP in numerical simulations. The LYZ carry information of the system that is not encoded in finite derivatives of the free energy. Our method thus serves as a procedure independent of the ordinary methods that rely on susceptibilities. We test the method for analyzing the CP in the three-dimensional three-state Potts model and show that the method can determine its location successfully with almost the same statistical error compared to the Binder-cumulant method. As byproducts, we also derive some useful properties of LYZ, especially their relation to the LYES, which play essential roles in controlling finite-size effects in numerical results.
@@
Ising model.— To illustrate the method, we start from a simple case of the conventional three-dimensional Ising (3D-Ising) model [3] described by the reduced temperature $t$ and external magnetic field $h$ on the cubic lattice of size $L^{3}$, having a CP at $(t,h)=(0,0)$ and a first-order phase transition at $h=0$ for $t<0$. We denote the partition function of this model as $Z(t,h,L^{-1})$, which satisfies $Z(t,h,L^{-1})=Z(t,-h,L^{-1})$. The LYZ of this model are the values of $h\in\mathbb{C}$ satisfying $Z(t,h,L^{-1})=0$ for a given $t\in\mathbb{R}$ [17, 18]. It is known as the Lee-Yang circle theorem that the LYZ distribute discretely on the pure-imaginary axis for finite $L$ [18]. In the following, we denote the LYZ with ${\rm Im}\ h>0$ as $h=h_{\rm LY}^{(n)}(t,L)$, where $n=1,2,\cdots$ labels different LYZ so that $0<{\rm Im}\ h_{\rm LY}^{(1)}(t,L)<{\rm Im}\ h_{\rm LY}^{(2)}(t,L)<\cdots$; the system has the other LYZ at $h=-h_{\rm LY}^{(n)}(t,L)$. Since $Z(t,h,L^{-1})$ at finite $L$ is a regular function of $t$ and $h$, $h_{\rm LY}^{(n)}(t,L)$ should also be regular functions of $t$ for finite $L$  [^1] The regularity of $h_{\rm LY}^{(n)}(t,L)$ is violated when $Z(t,h,L^{-1})=0$ has a multiple solution at some $t$ even for a regular $Z$. We, however, have numerically checked that such cases do not occur in the parameter range explored in this study..
@@
for sufficiently large $L$. In the 3D-Ising model, the values of the exponents, $y_{t}\simeq 1.588$ and $y_{h}\simeq 2.482$, have been analyzed with high precision [42, 15, 16]. Since the LYZ are given by zeros of Eq. (1), it immediately follows that the LYZ for different $L$ obey [43]
@@
for finite $n$. Since the right-hand side of Eq. (3) does not depend on $L$, only a possible asymptotic behavior of ${\tilde{h}}_{\rm LY}^{(n)}({\tilde{t}})$ for $\tilde{t}\to\infty$ is ${\tilde{h}}_{\rm LY}^{(n)}({\tilde{t}})\propto{\tilde{t}}^{y_{h}/y_{t}}$, which yields $h_{\rm LYES}(t)\propto t^{y_{h}/y_{t}}$ for $t>0$ [21].
@@
with $r_{nm}=X_{n}/X_{m}$ and $c_{nm}=r_{nm}(Y_{n}/X_{n}-Y_{m}/X_{m})$. Equation (6) shows that $R_{nm}(0,L)=r_{nm}$ is independent of $L$, while the slope at $t=0$ scales as $L^{y_{t}}$. In other words, $R_{nm}(t,L)$ for different $L$ intersects at $t=0$ as in Fig. 1 (a). This property is similar to the Binder cumulants [14] and would provide an alternative method to determine the critical temperature from the intersection point in numerical simulations.
@@
CP in general systems.— Next, we extend the argument to CPs in general systems that belong to the same universality class as the 3D-Ising model. We suppose a system whose partition function ${\cal Z}(\tau,\xi,l^{-1})$ is described by two variables $\tau$ and $\xi$ with $l$ being a dimensionless parameter proportional to the spatial size of the system. We also assume that this system has a first-order phase-transition line on the $\tau$$\xi$ plane that terminates at a CP at $(\tau,\xi)=(\tau_{\rm c},\xi_{\rm c})$ as depicted in Fig. 1 (b). In the following, we refer to zero points of the partition function for $\xi\in\mathbb{C}$ and $\tau\in\mathbb{R}$ as the LYZ, and denote the LYZ for ${\rm Im}\ \xi>0$ as $\xi=\xi_{\rm LY}^{(n)}(\tau,l)$, i.e. ${\cal Z}(\tau,\xi_{\rm LY}^{(n)}(\tau,l),l^{-1})=0$, where the definition of the label $n$ is the same as before.
Since there are only two relevant variables for the CP in the 3D-Ising universality class, the partition function in the vicinity of the CP for large $l$ is related to that of the Ising model as
@@
$l^{y_{h}}\check{h}\big{(}\tau,\xi_{\rm LY}^{(n)}(\tau,l)\big{)}=\tilde{h}_{\rm LY}^{(n)}\big{(}l^{y_{t}}\check{t}[\tau,\xi_{\rm LY}^{(n)}(\tau,l)]\big{)}.$ (10)
@@
$l^{y_{h}}\big{[}a_{21}\delta\tau+a_{22}(\xi_{\rm LY}^{(n)}(\tau,l)-\xi_{\rm c})\big{]}$
$=i\big{\{}X_{n}+Y_{n}l^{y_{t}}\big{[}a_{11}\delta\tau+a_{12}(\xi_{\rm LY}^{(n)}(\tau,l)-\xi_{\rm c})\big{]}\big{\}}+{\cal O}(\delta\tau^{2}),$ (11)
@@
Now, let us compare Eq. (15) with the Binder-cumulant method [14]. For locating a CP on the $\tau$$\xi$ plane, one may define the fourth-order Binder cumulant as ${\cal B}_{4}(\tau,l)={\rm min}_{\xi}[(\partial^{4}{\cal F}(\tau,\xi,l^{-1})/\partial\xi^{4})/(\partial^{2}{\cal F}(\tau,\xi,l^{-1})/\partial\xi^{2})^{2}]+3$ [45, 46, 47] with the free energy ${\cal F}(\tau,\xi,l^{-1})=-T\ln{\cal Z}(\tau,\xi,l^{-1})$ ) with temperature $T$. One then obtains [46, 48]
@@
$\frac{{\cal H}(\tau,\xi)}{T}=-\tau\sum_{\langle i,j\rangle}\delta_{\sigma_{i}\sigma_{j}}-\xi\sum_{i}\delta_{\sigma_{i},1},$ (17)
on the simple cubic lattice of size $L^{3}$ with periodic boundary conditions, where $\sigma_{i}$ takes three states $\sigma_{i}=1,2,3$ with the subscript denoting the lattice site and $\sum_{\langle i,j\rangle}$ represents the sum over all pairs of adjacent sites. As schematically shown in Fig. 1 (b), this model is $Z(3)$ symmetric and has a first-order phase transition at vanishing external field $\xi=0$, which is eventually terminated at a CP for $\xi>0$ that belongs to the 3D-Ising universality class [45]. In Ref. [45], this CP has been investigated by the Binder-cumulant method in relation to the CP in QCD with heavy-mass quarks [48, 47, 49].
@@
with $\tau\in\mathbb{R}$ and $\xi\in\mathbb{C}$ for each simulation parameter, where $\langle\cdot\rangle_{L}$ denotes the average over the configurations at $(\tau_{\rm sim},\xi_{\rm sim})$ of size $L^{3}$. The numerical cost to calculate Eq. (18) does not depend on $L$ and is negligibly small compared to that for updates of configurations. Whereas the analysis of Eq. (18) suffers from the overlapping problem when $(\tau,\xi)$ is largely deviated from $(\tau_{\rm sim},\xi_{\rm sim})$, we found that this problem is well suppressed in our analysis as demonstrated below.
@@
Table 1: Fit results of the CP parameters and $\chi^{2}/{\rm d.o.f}$.
fit data $\tau_{\rm c}$ $y_{t}$ $r_{n1}$ or $b_{4}$ $\chi^{2}/\rm{d.o.f}$
@@
To obtain the critical value $\tau=\tau_{\rm c}$ at the CP, we performed the four-parameter chi-square fits to the data of ${\cal R}_{nm}(\tau,L)$ for $L\geq 40$ ($12$ data points in total) with an ansatz ${\cal R}_{nm}(\tau,L)=r+c(\tau-\tau_{\rm c})L^{y_{t}}$ with $r$, $c$, $\tau_{\rm c}$, and $y_{t}$ being the fitting parameters. Effects of the second bracket in Eq. (15) are neglected since no deviation of the intersection point is visible for $L\geq 40$ in Fig. 2. The fit results are shown by the diamonds in Fig. 2 for $m=1$ and summarized in Table 1. The table shows that these results are consistent with each other, while smaller $m$ tends to give better statistics with fixed $n$.
@@
Summary and outlook.— In this Letter, we discussed the FSS of LYZ and showed that the intersection point of their ratios, Eq. (4) or Eq. (16), for various spatial volumes indicates a CP in general systems. This property can be used for numerical searches of CPs as an independent method from the conventional ones based on susceptibilities. Compared to the Binder-cumulant method [14], this method is advantageous in suppressing the finite-volume effects arising from $a_{12}\neq 0$. We applied the method to the numerical analysis of the CP in the three-state Potts model. Whereas we assumed the CP in the 3D-Ising universality class throughout the Letter for a simple presentation, this method, of course, can be extended to CPs in other universality classes.
In our numerical study, we limited our analysis to the LYZ with $n\leq 4$. However, one can use the LYZ for yet larger $n$, which will act to improve the statistics. The LYZ can also be used to determine $\xi_{\rm c}$ and the matrix $A$, where better control of the finite-volume effects and statistics will be realized by combined uses of Eqs. (13) and (14) for various $n$. The precise measurements of $r_{nm}$ in individual universality classes and the violation of the FSS (1) in specific models [15, 16], as well as the application of the method to CPs in various systems, such as the QCD critical point in lattice simulations, are other important future studies.
Acknowledgement.— We thank Shinji Ejiri for discussions in the early stage of this study. The authors also thank Koji Hukushima and Christian Schmidt for useful discussions. This work was supported in part by JSPS KAKENHI (Nos. JP19H05598, JP22K03593, JP22K03619, JP23H04507, JP24K07049), the Center for Gravitational Physics and Quantum Information (CGPQI) at Yukawa Institute for Theoretical Physics, the Research proposal-based use at the Cybermedia Center, Osaka University, and the Multidisciplinary Cooperative Research Program of the Center for Computational Sciences, University of Tsukuba.